OpenStudy (anonymous):

1.What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms? with work ? please

OpenStudy (anonymous):

262136 is the answer ?

OpenStudy (anonymous):

\[r=-2\] it looks like

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

No it's this The first term is 8, and the second is -16 Thus, this is a sum of the form (8)(-2)^n we have: Summation of 8(-2)^n from n = 0 to n = 14 = 8 [Summation of (-2)^n from n = 0 to n = 14] = 8 [((-2)^(15) - 1) / (-2 - 1)] = 87384

OpenStudy (anonymous):

so you have \[8(1+(-2)+(-2)^2+(-2)^3+...+(-2)^{14}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

end up with \[8\times \frac{(-2)^{15}-1}{-2-1}\] if a remember correctly

OpenStudy (anonymous):

what jackie said

OpenStudy (anonymous):

so the answer is 87384 ?

OpenStudy (anonymous):

yes i believe so.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem